1.what fraction of a bag of 24 marbles will each person have if it is shared among:
1.1.two people.
1.2.three people.
1.3.six people.
1.4.four people.
2.how many thirds make six wholes.
3.draw a shape of which one eighth of the whole has been coloured,and write the fraction of the shape that is uncoloured.
4.draw a shape of which seven-tenths of the whole is coloured and compare it with two-fifths of the whole.which one is bigger?illustrate your answer.
1
Expert's answer
2020-09-15T15:24:03-0400

1) 24 divided by the number of people, so:

1.1) 24:2 = 12

1.2) 24:3 = 8

1.3) 24:6 = 4

1.4) 24:4 = 6


2) The answer is 18, because:

"18:3 = 6"


3) fraction of the shape that is uncoloured: "\frac{7}{8}"


4)"\frac{7}{10}" (fist drawing)is bigger then "\frac{2}{5}" (second drawing)